OpenStudy (anonymous):

what's sin(45) = 8/X

OpenStudy (freckles):

sin(45)=sqrt(2)/2 solve this for X: \[\frac{\sqrt{2}}{2}=\frac{8}{X}\]

OpenStudy (anonymous):

huh?

OpenStudy (freckles):

your question is to solve for X: sin(45)=8/X right?

OpenStudy (anonymous):

correct

OpenStudy (freckles):

We know from the unit circle sin(45)=sqrt(2)/2

OpenStudy (freckles):

\[\frac{\sqrt{2}}{2}=\frac{8}{X}\] so we can replace sin(45) with sqrt(2)/2

OpenStudy (freckles):

now we solve this for X

OpenStudy (anonymous):

what is the answer out of these? a. 8sqrt 3 b. 8sqrt 2 c. 4 sqrt 2 d. 16 sqrt 2

OpenStudy (freckles):

you can cross multiplying

OpenStudy (freckles):

try it

OpenStudy (anonymous):

OMG Thanks!

OpenStudy (freckles):

remember to rationalize any unrationalized denominators